# Concerns and Values in Human-Robot Interactions: A Focus on Social Robotics

### v1

@@ 4.3.4 Value: Autonomy
With autonomy, we refer to the freedom of thinking, without being under the influence of external agents [11]. The concept differs from being physically autonomous from others, which we refer to as agency [72], as suggested by experts during the focus groups. During the scoping review, this value was initially linked to independence. However, the focus groups participants recommended separating it from independence because autonomy specifically refers to making decisions without external impositions. They emphasised that autonomy should focus on enhancing the user’s ability to make their own decisions, rather than on the robot’s capabilities. When a robot assists in decision-making, it should support the user’s best interests and enable them to accomplish tasks they otherwise could not.
@@
While robots can provide valuable companionship, there is a risk that users develop an emotional dependency on them. This could lead to isolation from other people [46], as highlighted during the focus groups. The user should not be allowed to form an excessive attachment to the robot when it impersonates a professional providing mental and emotional support. This reflects what happens between a patient and a human carer. However, attachment is still regarded as problematic in the case of a companion robot, even though the doctor-patient relationship does not play a role. Excessive attachment is thus always regarded as undesirable due to the artificial nature of the robot.
@@
Robots that interact with users can subtly influence their decisions and behaviour through persuasion [102] or nudging, as noted during the scoping review (e.g., [33]) and the focus groups. A common concern is that robots might promote actions, products, or behaviours that serve the interests of the company that produced them. This kind of manipulation undermines the user’s ability to make autonomous decisions, especially if the user is unaware of the robot’s influence, such as in the case of children or the elderly. In these situations, it is also important to assess the appropriate level of paternalism to apply [106]. Initially, this value was categorised under empowerment, but the experts suggested moving it under autonomy, as manipulation directly affects decision-making.
@@ 4.3.5 Value: Equity
Equity entails treating people differently based on the circumstances, to ensure an equal outcome [23]. In contrast, equality – treating everyone the same regardless of their situation – did not emerge as a relevant value during the focus groups discussions. Indeed, the experts suggested focusing on equity as a key value. They noted that equity is closely linked to demographics (who) and the environment (where), which can significantly impact a robot’s performance. For instance, the geographical location can influence how often a robot overheats and how easily it can be repaired.
@@
This topic represents the positive connotations of equity (e.g., [103, 19]). The robot treats all users fairly according to their individual needs and is sensitive to differences in physical ability, culture, and preferences. In addition, the experts suggested that robots could help reduce gender disparity, particularly in home settings, by taking on tasks like cleaning and cooking. This could ease the burden on women and promote a more balanced distribution of responsibilities.
@@
The focus groups participants and some publications (e.g., [8, 103]) voiced the concerns of certain communities and populations which felt not being part of the intended user base of certain robots. This can be caused by the physical aspects of the robot or by the interaction design. Especially when powered by data-driven technologies, the robots can reflect the bias present in the dataset on which their algorithms are trained, as highlighted by the focus groups participants. The experts also noted that a robot might struggle to accurately identify emotional states in children with varying cognitive abilities or socioeconomic backgrounds. Or, in another example, it could treat students in undesirable ways, such as reinforcing gender stereotypes.
@@
This topic extends the previous one on a small scale. While bias refers to a lack of fairness between different communities and cultures, exclusion instead underlines the importance of adapting the robot’s behaviours to the individual’s preferences, habits, economic background, etc., as emerged during the scoping review [105]. The experts noted that robots will inevitably make certain assumptions based on their programming about what constitutes “normal” behaviour. These assumptions could lead to exclusion and have problematic effects.
@@
As highlighted during the scoping review [86, 97], the users need to know who is responsible if a robot malfunctions or makes a harmful decision. Clear lines of accountability are essential so that those responsible can be held accountable for the robot’s actions. Establishing accountability helps build trust and ensures the fair and just deployment of robots. Originally, this topic was categorised under the value of truth, but focus groups participants suggested moving it under equity.
@@ 4.3.6 Value: Dignity
Dignity groups the topics relating to the respect owed to all humans and to the subject’s self-image [65]. Although it was not identified as an initial value during the scoping review, focus groups participants agreed that dignity should replace honour as a value, with agency considered separately.
@@
Respecting the dignity of users involves ensuring that the robots do not deceive or trick them into anthropomorphising their behaviour, as pointed out by the focus groups participants. For example, a robot that pretends to have emotions might lead users to form attachments or expectations that are not appropriate, leading to confusion or emotional harm. This is especially true in those segments of the population for whom the artificial nature of the robot may not be immediately evident [10].
@@
As emerged during the scoping review [62], interacting with a robot instead of a human being can sometimes be perceived as degrading, and this varies across cultures. For example, an elderly person might feel humiliated if a robot, rather than a human caregiver, is assigned to assist them with personal tasks. This can lead to a sense of loss of dignity, as the user might feel that they are being treated as less worthy of human care and attention.
@@
This represents the positive side of tricking, where a robot’s artificial nature is clearly disclosed to users. This means ensuring that users are aware that they are interacting with a machine and not a human, which helps to set appropriate expectations and avoid any misunderstandings as pointed out by the focus groups participants.
@@ 4.3.7 Value: Virtue
Repetitive interactions with others can in the long run shape our behaviour [38], and this applies to human-robot interaction as well. This value, suggested by the focus groups participants, addresses both the positive and negative influences a robot can have on its users.
@@
Robots have the potential to serve as mentors, setting good examples and fostering virtuous interactions [19], as pinpointed by the focus groups participants. For example, a robot designed to assist children with their education might encourage patience, perseverance, and kindness.
@@ 4.3.8 Value: Welfare
Finally, the topics that cover the positive influence of interacting with a robot were grouped under the value of welfare as identified during the scoping review. These span from the mental to the physical welfare of the user. Notably, while we found prescriptive instructions about how the robot should be made safe, we did not find many topics expressed in negative terms. For this reason, there are no topics with negative valence for this value. This might suggest that the capabilities of social robots are not considered advanced enough to pose a risk in this sense. However, we anticipate this could change in the future. Focus groups participants emphasised that welfare is crucial, particularly regarding usage constraints, such as ensuring a robot is not used by a child at night.
@@
Robots can play a significant role in education, not only by teaching users about various topics but also by educating them about thought processes and routines, as emerged from the scoping review (e.g., in [86, 51]). This educational role can be particularly valuable as a support to traditional learning methods, and not as an alternative.
@@
One of the unique advantages of robots is that they are not perceived as judgemental. For example, a kid might feel more comfortable discussing sensitive topics with a robot, knowing that the robot will not judge them. Being aware that the robot does not have sentiments lowers the stakes of the conversation, and can create a safe space for users to be honest and open [109].
@@
Safety is a fundamental concern in the design and use of robots, and this is not limited to interactions with vulnerable populations. Robots must be designed to be physically safe, meaning they should not pose any risk of injury to users, as also highlighted by our focus groups participants. This topic is present in previous literature with positive connotations [6], inviting us to reflect on this aspect during the design phase of a robot.
@@
This topic covers the mental aspects of welfare, while the previous covers the physical ones. Robots can contribute to users’ wellbeing by providing opportunities for disclosure, giving advice, and offering support in various aspects of life as highlighted during the scoping review [105]. For example, a robot designed to support mental health might provide suggestions for coping strategies, or simply offer a friendly ear.
@@ 4.4 Preliminary discussion and limitations
As a result of a scoping review and an elicitation process of expert perspectives, we identified a set of final values that are critical to guide HRI researchers in designing and deploying social robots in various contexts. The focus groups challenged and refined some of the values initially identified in the scoping review. For instance, honour was reconceptualized as dignity, emphasising respect for individual worth and autonomy. This refinement demonstrates the evolving understanding of values as they are contextualised in specific HRI scenarios showing the importance of participatory approaches. This phase presents different limitations. First, the values and topics identified may not exhaustively represent all perspectives in HRI. For example, the focus groups included experts from specific disciplines, and their views may differ from those of end-users, such as children, parents, or caregivers. Future work should involve these stakeholders to capture a more comprehensive perspective. Second, the scenarios used in the focus groups, while chosen in line with the focus of the scoping review, may have introduced biases by guiding participants towards limited contexts and issues. Future work should explore diverse scenarios that might reveal additional values not considered in this current work. Third, participants primarily represent Western perspectives. Future research should aim to include more diverse cultural contexts to ensure the generalizability of the findings. As a general observation, it was noted that contextualizing values within scenarios and prioritizing them is essential for designing robots that promote social good.
@@ 5 Phase 3: Design and Evaluation of the HRI-Value Compass
Building on the results so far presented, Phase 3 constitutes an effort to practically aid HRI researchers in identifying the values at play in HRI scenarios (RQ3). To this end, we developed the HRI Value Compass, which is available as a web application. This tool is designed to streamline the value elicitation process with a focus on the interaction between the end users and the robots. We evaluated the tool through a pilot study during which 20 HRI researchers used them to identify moral values in HRI scenarios – the same scenarios used during the focus groups presented in Section 4.
@@ 5.1 Design of the HRI value compass
The HRI Value Compass was designed following common principles of usability and user experience for websites [60]. The primary design goals for the tool were simplicity and ease of use. The tool is structured into three main sections: contextual information, values and topics, and results overview.
@@ 5.2 Evaluation of the HRI value compass
We invited participants to our online pilot study via email, providing them with a Qualtrics [^5] https://www.qualtrics.com/ link to an online questionnaire. In the online form, they signed an informed consent clause following the guidelines of Ghent University Ethics Committee and completed a demographic survey. In the first part, the participants read a description of an HRI scenario randomly selected from the set of four scenarios used during the focus groups activities. They then used the HRI Value Compass tool, answering questions about the context, values, and topics, and viewed the overview of the results. At any time, they had access to the full text of the scenario they were evaluating for reference. After evaluating the first scenario, participants repeated the same procedure a second time with a different scenario: every participant evaluated two scenarios. We limited the evaluation to two scenarios to prevent task overload.
In the second part, the participant completed the System Usability Scale (SUS) [15], the NASA Task Load Index (NASA-TLX) [76] in its raw variant, and a series of questions on the use of the HRI-Value Compass. The SUS questionnaire was administered to evaluate the usability of the HRI-Value Compass web tool, following a similar approach used in previous studies [90]. The NASA-TLX questionnaire was used to measure the cognitive workload experienced by participants while performing the task, as described by Gomer and Pagano [39]. The customised survey contained seven 5-point Likert scale questions reported in Table 5. Additionally, we included three open-ended questions to gather participants’ general opinions about the HRI-Value Compass.
@@ 5.3 Results: HRI value compass
The results of the SUS questionnaire were elaborated [15], obtaining a final score of 68.38, which scores slightly above the 50th percentile, in line with the average score of usability for digital tools [78]. The raw NASA-TLX returned an overall score of 48.17/100 (SD = 15.88), suggesting a somewhat high workload in completing the task mainly due to the mental demand (M = 69.75) and effort (M = 61.75) needed to analyse the scenarios and extracting the corresponding values. The average scores of the single scales are the following, with the standard deviation reported between parenthesis: mental demand 69.75 (25.10), physical demand 20.00 (24.33), temporal demand 33.75 (22.82), performance 60.50 (22.82), effort 61.75 (21.04), frustration 43.25 (30.96). The results of the custom Likert questionnaire are reported in Table 5.
@@
Suggested improvements include allowing the users to write a description of the use case and implementing a structured reporting format. One subject suggested including closed-ended questions, while another asked to clarify the separation between the topics’ definitions and the questions. We have not updated the tool following these indications yet, as we first seek to collect additional opinions on its improvement.
@@
Question Mean SD
I would recommend this tool to a colleague in HRI. 3.40 0.99
I think that this tool can be used also in group settings, to facilitate HRI design processes. 3.90 0.91
I found this tool very useful. 3.50 0.61
I will use this tool in the future in my HRI studies. 2.90 0.79
I found the references very useful. 3.50 1.00
The tool helps me to think about parts of the scenarios that I haven’t considered before. 4.10 1.07
I think that the tool helped me to extract values that I would not have thought about without using it. 4.05 1.00
@@ 5.4 Preliminary discussion and limitations
The main outcome of Phase 3 is the development of a web tool called HRI Value Compass, designed to assist and guide HRI researchers in identifying opportunities, concerns, and values in HRI. Our evaluation demonstrated that the tool is comprehensive, effective, and valuable for this purpose, with particular potential for use in collaborative settings. However, there is room for improvement, especially in enhancing its usability and reducing cognitive demands. The high cognitive workload scores may partly reflect participants’ unfamiliarity with the tool or the complexity of the scenarios analysed. Future studies could include a training phase to account for the learning curve and evaluate usability after repeated use. The study highlighted the tool’s potential for group use, but it did not explicitly test its performance in such settings. Future research should investigate how the tool performs in collaborative contexts to validate these findings. Although 46 participants were initially recruited, only 20 completed the study, potentially limiting the generalizability of the findings.
6 General Discussion and Conclusion
Findings from this work will assist HRI researchers in designing and deploying robotic systems that align with human values, societal norms, and ethical principles. First, we provided an overview of the current HRI literature to highlight the key values considered in HRI environments (Phase 1, addressing RQ1). Next, we defined a comprehensive list of topics and values with input from technology ethics experts (Phase 2, addressing RQ2). Finally, we developed and evaluated a web tool, “HRI Value Compass”, through a pilot study (Phase 3, addressing RQ3). This tool is designed to help HRI researchers practically identify and incorporate these values when designing robotic interactions.

### v2

@@ 4.3.4 Value: Autonomy
With autonomy, we refer to the freedom of thinking, without being under the influence of external agents [beauchamp2003methods]. The concept differs from being physically autonomous from others, which we refer to as agency [prunkl2022human], as suggested by experts during the focus groups. During the scoping review, this value was initially linked to independence. However, the focus group participants recommended separating it from independence because autonomy specifically refers to making decisions without external impositions. They emphasised that autonomy should focus on enhancing the user’s ability to make their own decisions, rather than on the robot’s capabilities. When a robot assists in decision-making, it should support the user’s best interests and enable them to accomplish tasks they otherwise could not.
@@
While robots can provide valuable companionship, there is a risk that users develop an emotional dependency on them. This could lead to isolation from other people [huber2016ethical], as highlighted during the focus groups. The user should not be allowed to form an excessive attachment to the robot when it impersonates a professional providing mental and emotional support. This reflects what happens between a patient and a human carer. However, attachment is still regarded as problematic in the case of a companion robot, even though the doctor-patient relationship does not play a role. Excessive attachment is thus always regarded as undesirable due to the artificial nature of the robot.
@@
Robots that interact with users can subtly influence their decisions and behaviour through persuasion [vollmer2018children] or nudging, as noted during the scoping review (e.g., [fiske2019your]) and the focus groups. A common concern is that robots might promote actions, products, or behaviours that serve the interests of the company that produced them. This kind of manipulation undermines the user’s ability to make autonomous decisions, especially if the user is unaware of the robot’s influence, such as in the case of children or the elderly. In these situations, it is also important to assess the appropriate level of paternalism to apply [white2013manipulation]. Initially, this value was categorised under empowerment, but the experts suggested moving it under autonomy, as manipulation directly affects decision-making.
@@ 4.3.5 Value: Equity
Equity entails treating people differently based on the circumstances, to ensure an equal outcome [cook1983distributive]. In contrast, equality – treating everyone the same regardless of their situation – did not emerge as a relevant value during the focus group discussions. Indeed, the experts suggested focusing on equity as a key value. They noted that equity is closely linked to demographics (who) and the environment (where), which can significantly impact a robot’s performance. For instance, the geographical location can influence how often a robot overheats and how easily it can be repaired.
@@
This topic represents the positive connotations of equity (e.g., [wagman2021beyond, cappuccio2021can]). The robot treats all users fairly according to their individual needs and is sensitive to differences in physical ability, culture, and preferences. In addition, the experts suggested that robots could help reduce gender disparity, particularly in home settings, by taking on tasks like cleaning and cooking. This could ease the burden on women and promote a more balanced distribution of responsibilities.
@@
The focus group participants and some publications (e.g., [baisch2023elders, wagman2021beyond]) voiced the concerns of certain communities and populations who felt not part of the intended user base of certain robots. This can be caused by the physical aspects of the robot or by the interaction design. Especially when powered by data-driven technologies, the robots can reflect the bias present in the dataset on which their algorithms are trained, as highlighted by the focus group participants. The experts also noted that a robot might struggle to accurately identify emotional states in children with varying cognitive abilities or socioeconomic backgrounds. Or, in another example, it could treat students in undesirable ways, such as reinforcing gender stereotypes.
@@
This topic extends the previous one on a small scale. While bias refers to a lack of fairness between different communities and cultures, exclusion instead underlines the importance of adapting the robot’s behaviours to the individual’s preferences, habits, economic background, etc., as emerged during the scoping review [wheeler2023confucianism]. The experts noted that robots will inevitably make certain assumptions based on their programming about what constitutes “normal” behaviour. These assumptions could lead to exclusion and have problematic effects.
@@
As highlighted during the scoping review [smakman2020acceptable, van2020teachers], the users need to know who is responsible if a robot malfunctions or makes a harmful decision. Clear lines of accountability are essential so that those responsible can be held accountable for the robot’s actions. Establishing accountability helps build trust and ensures the fair and just deployment of robots. Originally, this topic was categorised under the value of truth, but focus group participants suggested moving it under equity.
@@ 4.3.6 Value: Dignity
Dignity groups the topics relating to the respect owed to all humans and to the subject’s self-image [nordenfelt2004varieties]. Although it was not identified as an initial value during the scoping review, focus group participants agreed that dignity should replace honour as a value, with agency considered separately.
@@
Respecting the dignity of users involves ensuring that the robots do not deceive or trick them into anthropomorphising their behaviour, as pointed out by the focus group participants. For example, a robot that pretends to have emotions might lead users to form attachments or expectations that are not appropriate, leading to confusion or emotional harm. This is especially true in those segments of the population for whom the artificial nature of the robot may not be immediately evident [battistuzzi2018embedding].
@@
As emerged during the scoping review [misselhorn2013ethical], interacting with a robot instead of a human being can sometimes be perceived as degrading, and this varies across cultures. For example, an elderly person might feel humiliated if a robot, rather than a human caregiver, is assigned to assist them with personal tasks. This can lead to a sense of loss of dignity, as the user might feel that they are being treated as less worthy of human care and attention.
@@
This represents the positive side of tricking, where a robot’s artificial nature is clearly disclosed to users. This means ensuring that users are aware that they are interacting with a machine and not a human, which helps to set appropriate expectations and avoid any misunderstandings as pointed out by the focus group participants.
@@ 4.3.7 Value: Virtue
Repetitive interactions with others can in the long run shape our behaviour [gardner2020does], and this applies to human-robot interaction as well. This value, suggested by the focus group participants, addresses both the positive and negative influences a robot can have on its users.
@@
Robots have the potential to serve as mentors, setting good examples and fostering virtuous interactions [cappuccio2021can], as pinpointed by the focus group participants. For example, a robot designed to assist children with their education might encourage patience, perseverance, and kindness.
@@ 4.3.8 Value: Welfare
Finally, the topics that cover the positive influence of interacting with a robot were grouped under the value of welfare as identified during the scoping review. These span from the mental to the physical welfare of the user. Notably, while we found prescriptive instructions about how the robot should be made safe, we did not find many topics expressed in negative terms. For this reason, there are no topics with negative valence for this value. This might suggest that the capabilities of social robots are not considered advanced enough to pose a risk in this sense. However, we anticipate this could change in the future. Focus group participants emphasised that welfare is crucial, particularly regarding usage constraints, such as ensuring a robot is not used by a child at night.
@@
Robots can play a significant role in education, not only by teaching users about various topics but also by educating them about thought processes and routines, as emerged from the scoping review (e.g., in [smakman2020acceptable, jung2023enjoy]). This educational role can be particularly valuable as a support to traditional learning methods, and not as an alternative.
@@
One of the unique advantages of robots is that they are not perceived as judgmental. For example, a kid might feel more comfortable discussing sensitive topics with a robot, knowing that the robot will not judge them. Being aware that the robot does not have sentiments lowers the stakes of the conversation, and can create a safe space for users to be honest and open [yeung2023living].
@@
Safety is a fundamental concern in the design and use of robots, and this is not limited to interactions with vulnerable populations. Robots must be designed to be physically safe, meaning they should not pose any risk of injury to users, as also highlighted by our focus group participants. This topic is present in previous literature with positive connotations [amirabdollahian2013accompany], inviting us to reflect on this aspect during the design phase of a robot.
@@
This topic covers the mental aspects of welfare, while the previous covers the physical ones. Robots can contribute to users’ wellbeing by providing opportunities for disclosure, giving advice, and offering support in various aspects of life as highlighted during the scoping review [wheeler2023confucianism]. For example, a robot designed to support mental health might provide suggestions for coping strategies, or simply offer a friendly ear.
@@ 4.4 Preliminary discussion and limitations
As a result of a scoping review and an elicitation process of expert perspectives, we identified a set of final values that are critical to guide HRI researchers in designing and deploying social robots in various contexts. The focus groups challenged and refined some of the values initially identified in the scoping review. For instance, honour was reconceptualised as dignity, emphasising respect for individual worth and autonomy. This refinement demonstrates the evolving understanding of values as they are contextualised in specific HRI scenarios showing the importance of participatory approaches. This phase presents different limitations. First, the values and topics identified may not exhaustively represent all perspectives in HRI. For example, the focus groups included experts from specific disciplines, and their views may differ from those of end-users, such as children, parents, or caregivers. Future work should involve these stakeholders to capture a more comprehensive perspective. Second, the scenarios used in the focus groups, while chosen in line with the focus of the scoping review, may have introduced biases by guiding participants towards limited contexts and issues. Future work should explore diverse scenarios that might reveal additional values not considered in this current work. Third, participants primarily represent Western perspectives. Future research should aim to include more diverse cultural contexts to ensure the generalizability of the findings. As a general observation, it was noted that contextualising values within scenarios and prioritising them is essential for designing robots that promote social good.
@@ 5 Phase 3: Design and Evaluation of the HRI-Value Compass
Building on the results so far presented, Phase 3 constitutes an effort to practically aid HRI researchers in identifying the values at play in HRI scenarios (RQ3). To this end, we developed the HRI Value Compass, which is available as a web application. This tool is designed to streamline the value elicitation process with a focus on the interaction between the end users and the robots. We evaluated the tool through a pilot study during which 20 HRI researchers used it to identify moral values in HRI scenarios – the same scenarios used during the focus groups presented in Section 4.
@@ 5.1 Design of the HRI Value Compass
The HRI Value Compass was designed following common principles of usability and user experience for websites [matera2006web]. The primary design goals for the tool were simplicity and ease of use. The tool is structured into three main sections: contextual information, values and topics, and results overview.
@@ 5.2 Evaluation of the HRI Value Compass
We invited participants to our online pilot study via email, providing them with a Qualtrics [^5] https://www.qualtrics.com/ link to an online questionnaire. In the online form, they signed an informed consent clause following the guidelines of Ghent University Ethics Committee and completed a demographic survey. In the first part, the participants read a description of an HRI scenario randomly selected from the set of four scenarios used during the focus group activities. They then used the HRI Value Compass tool, answering questions about the context, values, and topics, and viewed the overview of the results. At any time, they had access to the full text of the scenario they were evaluating for reference. After evaluating the first scenario, participants repeated the same procedure a second time with a different scenario: every participant evaluated two scenarios. We limited the evaluation to two scenarios to prevent task overload.
In the second part, the participant completed the System Usability Scale (SUS) [brooke2013sus], the NASA Task Load Index (NASA-TLX) [rubio2004evaluation] in its raw variant, and a series of questions on the use of the HRI-Value Compass. The SUS questionnaire was administered to evaluate the usability of the HRI-Value Compass web tool, following a similar approach used in previous studies [spitale2021composing]. The NASA-TLX questionnaire was used to measure the cognitive workload experienced by participants while performing the task, as described by gomer2011nasa. The customised survey contained seven 5-point Likert scale questions reported in Table 6. Additionally, we included three open-ended questions to gather participants’ general opinions about the HRI-Value Compass.
@@ 5.3 Results: HRI Value Compass
The results of the SUS questionnaire were elaborated [brooke2013sus], obtaining a final score of 68.38, which scores slightly above the 50th percentile, in line with the average score of usability for digital tools [sauro2016quantifying]. The raw NASA-TLX returned an overall score of 48.17/100 (SD = 15.88), suggesting a somewhat high workload in completing the task mainly due to the mental demand (M = 69.75) and effort (M = 61.75) needed to analyse the scenarios and extract the corresponding values. The average scores of the single scales are the following, with the standard deviation reported between parentheses: mental demand 69.75 (25.10), physical demand 20.00 (24.33), temporal demand 33.75 (22.82), performance 60.50 (22.82), effort 61.75 (21.04), frustration 43.25 (30.96). The results of the custom Likert questionnaire are reported in Table 6.
@@
Suggested improvements include allowing users to write a description of the use case and implementing a structured reporting format. One subject suggested including closed-ended questions, while another asked to clarify the separation between the topics’ definitions and the questions. We have not updated the tool following these indications yet, as we first seek to collect additional opinions on its improvement.
@@
Question Mean SD I would recommend this tool to a colleague in HRI. 3.40 0.99 I think that this tool can be used also in group settings, to facilitate HRI design processes. 3.90 0.91 I found this tool very useful. 3.50 0.61 I will use this tool in the future in my HRI studies. 2.90 0.79 I found the references very useful. 3.50 1.00 The tool helps me to think about parts of the scenarios that I haven’t considered before. 4.10 1.07 I think that the tool helped me to extract values that I would not have thought about without using it. 4.05 1.00
@@ 5.4 Preliminary discussion and limitations
The main outcome of Phase 3 is the development of a web tool called HRI Value Compass, designed to assist and guide HRI researchers in identifying opportunities, concerns, and values in HRI. Our evaluation demonstrated that the tool is comprehensive, effective, and valuable for this purpose, with particular potential for use in collaborative settings. However, there is room for improvement, especially in enhancing its usability and reducing cognitive demands. The high cognitive workload scores may partly reflect participants’ unfamiliarity with the tool or the complexity of the scenarios analysed. Future studies could include a training phase to account for the learning curve and evaluate usability after repeated use. The study highlighted the tool’s potential for group use, but it did not explicitly test its performance in such settings. Future research should investigate how the tool performs in collaborative contexts to validate these findings. Although 46 participants were initially recruited, only 20 completed the study, potentially limiting the generalizability of the findings. Furthermore, the recruitment through professional connections, which was necessary as the participants had to be HRI researchers, might have influenced the results through positive bias.
6 General Discussion